OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=−8csc(5x+π/8)−2. smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=−8csc(5x+π/8)−2. smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

hey i'm back sooner than i thought; did i make a mistake? you are posting the last question i answered

OpenStudy (anonymous):

none of the answer work

OpenStudy (anonymous):

i donno why

OpenStudy (anonymous):

on the one you reposted?

OpenStudy (anonymous):

this one at the top of the page

OpenStudy (anonymous):

sorry, I found my mistake post following

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

y=−8csc(5x+π/8)−2 I missed the NEGATIVE in front of the 8 (plus another error) period=2pi/5 half period = pi/5 = 8pi/40 ANSWER ORDER CHANGED first negative asymptote: x= -pi/40 (shift left) smallest non-negative asymptote: x= pi/40 + pi/5 = 7pi/40 second non-negative asymptote: x= 7pi/40 + pi/5 = 15pi/40 = 3pi/8

OpenStudy (anonymous):

yea they worked thx for the help this is the last one if you dont mind Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=10tan(5x−π/10). smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

y=10tan(5x−π/10) period = pi/5; half period = pi/10 Inflection point at x=pi/50 smallest non-negative asymptote: x= pi/50 + pi/10 = 6pi/50 = 3pi/25 second non-negative asymptote: x= 3pi/25 + 5pi/25 = 8pi/25 first negative asymptote: x= 3pi/25 - 5pi/25 = -2pi/25

OpenStudy (anonymous):

wow thanks very much for all the help

OpenStudy (anonymous):

ur welcome :})